OpenStudy (anonymous):

Solve for v. -8v = -7v + 9

TheSmartOne (thesmartone):

Add 7v on both sides :)

TheSmartOne (thesmartone):

-8v + 7v = 9 v(-8+7) = 9 What is 7 - 8 =? Remember, 7 - 8 = -8 + 7

OpenStudy (anonymous):

Okay.

TheSmartOne (thesmartone):

...